OpenStudy (anonymous):

find an equation the goes trough (1,8) and (-2,-1)

OpenStudy (anonymous):

y-y0=m(x-x0)

OpenStudy (anonymous):

8--1= 9 1--2= 3 m= 9/3 so the equation would be F(x)= 9/3x + b -----------------> plug (x,y) = (1,8) -----------------> you get --> 8 = 9/3*1 + b -----------------> b = 8*3/9 so the equation is F(x)= 9/3x + 8*3/9